i must find the exact value of six trig functions of pi/4 radians.
sin pi/4 csc pi/4
cos pi/4 sec pi/4
tan pi/4 cot pi/4
please help me i have no idea how to do this. thank you i really appreciate any help.

Found 2 solutions by stanbon, CharlesG2:
Answer by stanbon(75887)

find the exact value of six trig functions of pi/4 radians.
Draw a 45/45/90 right triangle
Let the hypotenuse be 1.
Then, using Pytagoras, each side is sqrt(2)
----
sin (pi/4) = 1/sqrt(2)
csc pi/4 = sqrt(2)/1 = sqrt(2)
--------------------------------
cos (pi/4) = 1/sqrt(2)
sec (pi/4) = sqrt(2)
--------------------------------
tan (pi/4)= sqrt(2)/sqrt(2) = 1
cot (pi/4)= sqrt(2)/sqrt(2) = 1
==============================
Cheers,
Stan H.
==============================

Answer by CharlesG2(834)

I must find the exact value of six trig functions of pi/4 radians.
sin pi/4 csc pi/4
cos pi/4 sec pi/4
tan pi/4 cot pi/4
please help me i have no idea how to do this. thank you i really appreciate any help.
1. have your calculating device set to radians
2. sin = 1/csc; cos = 1/sec; tan = sin/cos = 1/cot
sin pi/4 = sqrt(2)/2
cos pi/4 = sqrt(2)/2
tan pi/4 = 1
3. the other 3 are the reciprocals